IF I HAD 87 CALLS COME IN, AND OUT OF THOSE I RETURNED 66 CALLS, WHAT PERCENTAGE WOULD THAT 66 CALLS EQUAL?
Answer by nerdybill(7384)

IF I HAD 87 CALLS COME IN, AND OUT OF THOSE I RETURNED 66 CALLS, WHAT PERCENTAGE WOULD THAT 66 CALLS EQUAL?
66/87 * 100 = 0.759 * 100 = 75.9%